# Interpolating Video-LLMs: Toward Longer-sequence LMMs in a Training-free Manner

### v1

@@ 3.2 Alternative Video Token Rearrangement
To address challenge , we enable the video encoder and alignment projector to map video input with an increased number of frames into the text feature space. As shown in Fig.2, the pre-trained video encoder is configured to process a fixed-length video sequence, $\mathbf{X}_{v}$, consisting of $N$ frames. The most straightforward approach is repeatedly using the encoder and projector to process $m\cdot N$ frames by sampling and encoding multiple subsets of frames $m$ times. Once we have collected $m$ groups of video tokens, we can concatenate them together to get the whole long video’s visual tokens.
However, since these video encoders and projectors are pre-trained to process only $N$ frames collectively, such repetitive usages lead to distorted temporal representations in the video token sequence. The concatenating video tokens may carry the frame’s relative information—such as the inconsistency between the tokens of the $N$-th frame and those of the $(N+1)$-th frame (i.e., tokens of the $1$-th frame and those of the $(N+1)$-th frame share more similarity). This discrepancy emerges because the transformer-processed tokens are not designed to naturally accommodate the increased spatial and temporal variety, potentially causing misalignments and inconsistencies in the encoded video features [Xiao et al., 2023, Shang et al., 2024a].
@@
To overcome this issue, we propose an alternative video token rearrangement technique that preserves the temporal consistency of the video tokens. The technique allows the generation of a infinite number of video tokens, which enable model to process extended video sequences without requiring retraining. The module employs a dynamic rearrangement of input frames and their corresponding tokens. Here’s how the rearrangement works:
@@
    To enhance the temporal coverage, the number of frames sampled is increased to $m\cdot N$, and the video is divided into subsequences to capture various temporal segments. For example: The first subsequence might include Frames {#$\{\#1,\#(m+1),\cdots\#(mN-m+1)\}$, providing some new but temporally overlapped “view” of these segments. Similarly, the rest $i$-th subsequence might consist of Frames {#$\{\#i,\#(m+i),\cdots\#(mN-m+i)\}$. These subsequences are processed separately through the same fixed video encoder and projector: The $i$-th subsequence processes to yield visual tokens $\mathbf{Z}_{v,i}$ and corresponding projected tokens $\mathbf{H}_{v,i}$. This approach increases the amount of visual data that can be processed without modifying the architecture or requiring retraining of the encoder and projector.
@@ 3.3 Interpolating Video-LLM backbone
Addressing the challenge , we introduce a post-training interpolation method for the Video-LLM’s backbone, specifically targeting its positional embedding. As discussed in the related work (Sec.2), Rotary Position Embedding (RoPE) [Su et al., 2024] is a widely-used positional encoding scheme adopted by several prominent LLMs, including Llama, Phi, Mistral, and Gemma. Consequently, RoPE has become the standard method for position embedding in Video-LLMs.
RoPE Fundamentals: RoPE integrates positional information into both the query vector $\mathbf{q}$ and key vector $\mathbf{k}$ of transformer models. Specifically, for a sequence of tokens $s_{1},s_{2},\cdots,s_{L}$ with embeddings $\mathbf{x}_{1},\cdots,\mathbf{x}_{L}\in\mathbb{R}^{D}$ ($D$ is the dimension of the embedding), RoPE applies the following transformations:
@@
where $\text{Re}(\mathbf{q},\mathbf{k})$ is the real part of the inner product of $\mathbf{q}$, $\mathbf{k}$, and $g(\cdot)$ is an abstract mapping function. In real coordinates, the RoPE can be expressed as follows:
@@
Context Window Scaling: Given the fixed context lengths of pre-trained Video-LLMs, a significant question is how to extend these lengths cost-effectively. Utilizing the inherent flexibility of RoPE, we modify the embedding functions to handle extended sequences without extensive retraining:
@@
NTK-aware Interpolation: The “NTK-aware” interpolation further refines this adjustment by recalibrating the base of the RoPE function:
@@ 3.4 Efficiency Bottleneck Analysis and a Post-training Solution
Inference Cost Analysis. To quantify the inference costs, we employ a modified roofline-based LLM-Viewer analysis, initially developed in [Yuan et al., 2024]. Consider a typical scenario in which each frame is represented by 256 visual tokens. In standard Video-LLaVA [Lin et al., 2023] configurations, 8 frames represent a video, totaling 2048 visual tokens. With the implementation of INTP, more frames can be processed, escalating the number of visual tokens significantly. This increase is thoroughly analyzed in Tab. 1, which details the computational cost implications for the LLM backbone’s decoding process, especially focusing on the model latency and memory cost.
@@
To address this issue, we introduce a KV-cache compression technique aimed at optimizing KV-cache storage efficiency. Post-training quantization (PTQ) is utilized to convert full-precision tensor inputs into quantized tensors, focusing primarily on selecting optimal quantization parameters. The process requires only two parameters: the scaling factor $S$ and the zero point $Z$, which are crucial for the quantization process [Shang et al., 2024b, Yuan et al., 2024]. To formalize, once appropriate values for $S$ and $Z$ are determined, a full-precision key (or value) tensor can be quantized as follows:
@@
where $\left[p_{min},p_{max}\right]$ is the quantization range determined by bit-width, for 2 bit integer, the bins are $\{-2,-1,0,1\}$. However, direct application of this method to KV quantization is not straightforward because KV tensors require input to be accessed. To address this, we employ a calibration dataset (a significantly smaller set of input samples compared to the training dataset) to collect the necessary tensors. Once the full-precision KV tensors at the $k$-th layer, $\mathbf{K}^{k}_{F}$ and $\mathbf{V}^{k}_{F}$, are obtained, they can be quantized using the same method to produce $\mathbf{K}^{k}_{Q}$ and $\mathbf{V}^{k}_{Q}$ based on the quantization equation above.
@@ 4 Experiments
We detail the experimental setup and model configurations in Sec. 4.1, followed by an analysis of the quantitative and qualitative performance of our approach in Sec. 4.2 and Sec. 4.4. Finally, we demonstrate the effectiveness of each component in our model in Sec.4.3.
@@ 4.1 Experimental Setup
Model Settings. We use the Video-LLaVA [Lin et al., 2023] framework to explore Video-LLM interpolation [^1] We develop our method based on the Video-LLaVA codebase: https://github.com/PKU-YuanGroup/Video-LLaVA. The evaluation of our approach is conducted using the IG-VLM [Kim et al., 2024] codebase: https://github.com/imagegridworth/IG-VLM.. We employ the Vicuna-7B v1.5 as the LLM backbone, mirroring the architecture used in Video-LLaVA. The visual encoders are adapted from LanguageBind, and the text tokenization is handled by the LLaMA tokenizer, which includes approximately 32,000 classes. The projection layers consist of 2 fully connected layers, facilitating the integration of visual and textual data.
Data and Training Details. Consistent with our focus on developing a training-free method, our approach requires no traditional data training processes. This aspect underscores the innovative nature of our interpolation method, which leverages existing model architectures and tools without the need for retraining or additional data, simplifying the implementation and reducing the computational overhead typically associated with training new models.
Datasets and Evaluation Metric. We assess INTP across a variety of zero-shot video question-answer benchmarks, which we classify as either open-ended or multiple-choice, depending on the type of question posed [Kim et al., 2024]. Our evaluation for open-ended VQA includes benchmarks like MSVD-QA [Xu et al., 2017], MSRVTT-QA [Xu et al., 2017], and ActivityNet-QA [Yu et al., 2019]. We utilize GPT-assisted assessments as outlined in Video-ChatGPT [Maaz et al., 2023], which provide a thorough evaluation of the model’s response accuracy and correctness. For multiple-choice VQA tasks, we evaluate the model’s performance on NExT-QA [Xiao et al., 2021] and EgoSchema [Mangalam et al., 2024]. We determine accuracy by the model’s ability to correctly select the appropriate answer from the provided options.
Table 2: A comparison of different LVMs on video reasoning benchmarks. Like Video-LLaVA [Lin et al., 2023], ChatGPT-Assistant is employed to evaluate the following performance. The version of ChatGPT is “gpt-3.5-turbo”.
@@
FrozenBiLM 1B 32.2 - 16.8 - 24.7 -
VideoChat 7B 56.3 2.8 45.0 2.5 - 2.2
LLaMA-Adapter 7B 54.9 3.1 43.8 2.7 34.2 2.7
Video-LLaMA 7B 51.6 2.5 29.6 1.8 12.4 1.1
Video-ChatGPT 7B 64.9 3.3 49.3 2.8 35.2 2.7
Video-LLaVA 7B 70.7 3.9 59.2 3.5 45.3 3.3
@@
FrozenBiLM 1B 26.9 - - - -
InternVideo 1.3B 32.1 48.0 43.4 65.1 59.1
Sevilla 7B - 61.3 61.5 75.6 63.6
Video-ChatGPT 7B - 61.9 57.4 69.9 61.7
Video-LLaVA 7B 37.0 61.2 54.2 71.1 60.5
@@ 4.4 Qualitative Results
In addition to the quantitative results, we visualize the question-and-answer samples within the ActivityNet dataset to assess the qualitative performance of our approach in Fig.3. These visualizations demonstrate that the INTP-enhanced Video-LLM is capable of effectively handling longer sequences of frames, showcasing a tangible improvement in processing and understanding extended video content. This qualitative assessment further validates the robustness of INTP-Video-LLM in real-world scenarios.
Figure 3: Video content and questions are from the ActivityNet benchmark [Yu et al., 2019]. The standard Video-LLaVA model exhibits limitations in accurately answering the benchmark’s questions, primarily due to its inability to process a sufficient number of frames (only the pink labeled frames are fed into Video-LLaVA). This constraint significantly hinders its effectiveness in managing complex video question-answering tasks. Nevertheless, with the implementation of INTP, the enhanced Video-LLM is capable of processing extended sequences of video frames. This improvement not only addresses the frame limitations but also substantially enhances the model’s understanding in complex video question-answering scenarios.
@@ 5 Conclusion
In conclusion, this paper presents INTP, an innovative interpolation method for Video-LLMs that effectively addresses the limitations of current models in processing extended video sequences. By developing an alternative video token rearrangement technique and a training-free LLM context window extension, INTP enables Video-LLMs to handle significantly more visual data (32 frames) without the need for additional training or substantial computational resources. Furthermore, the introduction of a KV-cache compression module optimizes memory usage during inference, enhancing deployment efficiency. These advancements not only extend the practical utility of Video-LLMs but also demonstrate the potential of training-free approaches in advancing Video-LLMs.
Societal Impacts. In this study, we introduce INTP, a interpolation method designed to enhance the capabilities of Video-LLMs for processing extended video sequences without requiring additional training. This approach can democratize the access to advanced video processing technologies by reducing computational demands. Despite these advancements, it is important to note that INTP does not address potential security concerns related to the misuse of Video-LLMs. The ease of accessibility and enhanced capabilities could potentially be exploited by malicious actors to generate or manipulate video content in harmful ways. Thus, while INTP expands the practical applications of Video-LLMs and reduces barriers to their use, it also necessitates careful consideration of the ethical implications and security measures associated with their deployment.
6 Acknowledgement
We thanks for the help from Yang Zhou, Haoxuan Wang, Zhihang Yuan for discussing the long-content LLM and KV-cache compression.

### v2

@@ 3.2 Alternative Video Token Rearrangement
To address challenge , we enable the video encoder and alignment projector to map video inputs with an increased number of frames into the text feature space. As shown in Fig. 2, the pre-trained video encoder is configured to process a fixed-length video sequence, $\mathbf{X}_{v}$, consisting of $N$ frames. The most straightforward approach is repeatedly using the encoder and projector to process $m\cdot N$ frames by sampling and encoding multiple subsets of frames $m$ times. Once we have collected $m$ groups of video tokens, we can concatenate them together to get the whole long video’s visual tokens.
However, since these video encoders and projectors are pre-trained to process only $N$ frames collectively, such repetitive usages lead to distorted temporal representations in the video token sequence. The concatenating video tokens may carry the frame’s relative information—such as the inconsistency between the tokens of the $N$-th frame and those of the $(N+1)$-th frame (i.e., tokens of the $1$-th frame and those of the $(N+1)$-th frame end up sharing more similarity). This discrepancy emerges because the transformer-processed tokens are not designed to naturally accommodate the increased spatial and temporal variety, potentially causing misalignments and inconsistencies in the encoded video features [Xiao et al., 2023, Shang et al., 2024a].
@@
To overcome this issue, we propose a video token rearrangement technique that preserves the temporal consistency of the video tokens. The technique allows the generation of an increased number of temporal-consistent video tokens, which enable the model to process extended video sequences without requiring retraining. The module employs a dynamic rearrangement of input frames and their corresponding tokens. Here’s how the rearrangement works:
@@
    To enhance temporal coverage, the number of frames sampled is increased to $m\cdot N$, and the video is divided into subsequences to capture various temporal segments. For example: The first subsequence might include Frames {#$\{\#1,\#(m+1),\cdots\#(mN-m+1)\}$, providing some new but temporally overlapped “view” of these segments. Similarly, the rest $i$-th subsequence might consist of Frames {#$\{\#i,\#(m+i),\cdots\#(mN-m+i)\}$. These subsequences are processed separately through the same fixed video encoder and projector: The $i$-th subsequence processes to yield visual tokens $\mathbf{Z}_{v,i}$ and corresponding projected tokens $\mathbf{H}_{v,i}$. This approach increases the amount of visual data that can be processed without modifying the architecture or requiring retraining of the encoder and projector.
@@ 3.3 Interpolating Video-LLM backbone
To address challenge , we introduce a post-training interpolation method for the Video-LLM’s backbone, specifically targeting its positional embedding. As discussed in the related work (Sec. 2), Rotary Position Embedding (RoPE) [Su et al., 2024] is a widely-used positional encoding scheme adopted by several prominent LLMs, including Llama, Phi, Mistral, and Gemma. Consequently, RoPE has become the standard method for position embedding in Video-LLMs.
3.3.1 RoPE Fundamentals
RoPE integrates positional information into both the query vector $\mathbf{q}$ and key vector $\mathbf{k}$ of transformer models. Specifically, for a sequence of tokens $s_{1},s_{2},\cdots,s_{L}$ with embeddings $\mathbf{x}_{1},\cdots,\mathbf{x}_{L}\in\mathbb{R}^{D}$ ($D$ is the dimension of the embedding), RoPE applies the following transformations:
@@ 3.3.1 RoPE Fundamentals
where is the conjugate of a complex number, $\text{Re}(\mathbf{q},\mathbf{k})$ is the real part of the inner product of $\mathbf{q}$, $\mathbf{k}$, and $g(\cdot)$ is an abstract mapping function. In real coordinates, the RoPE can be expressed as follows:
@@
3.3.2 Context Window Scaling
Given the fixed context lengths of pre-trained Video-LLMs, a significant question is how to extend these lengths cost-effectively. Utilizing the inherent flexibility of RoPE, we modify the embedding functions to handle extended sequences without extensive retraining:
@@ 3.3.2 Context Window Scaling
3.3.3 NTK-aware Interpolation
The “NTK-aware” interpolation further refines this adjustment by recalibrating the base of the RoPE function:
@@ 3.4 Efficiency Bottleneck Analysis and a Post-training Solution
3.4.1 Inference Cost Analysis
To quantify the inference costs, we employ a modified roofline-based LLM-Viewer analysis, initially developed in [Yuan et al., 2024]. Consider a typical scenario in which each frame is represented by 256 visual tokens. In standard Video-LLaVA [Lin et al., 2024] configurations, 8 frames represent a video, totaling 2048 visual tokens. With the implementation of INTP, more frames can be processed, escalating the number of visual tokens significantly. This increase is thoroughly analyzed in Tab. 1, which details the computational cost implications for the LLM backbone’s decoding process, especially focusing on the model latency and memory cost.
@@ 3.4.1 Inference Cost Analysis
To address this issue, we introduce a KV-cache compression technique aimed at optimizing KV-cache storage efficiency. Post-training quantization (PTQ) is utilized to convert full-precision tensor inputs into quantized tensors, focusing primarily on selecting optimal quantization parameters. The process requires only two parameters: the scaling factor $S$ and the zero point $Z$, which are crucial for the quantization process [Shang et al., 2024b, Yuan et al., 2024]. Formally, once appropriate values for $S$ and $Z$ are determined, a full-precision key (or value) tensor can be quantized as follows:
@@
where $\left[p_{min},p_{max}\right]$ is the quantization range determined by bit-width, for 2 bit integer, and the bins are $\{-2,-1,0,1\}$. However, direct application of this method to KV quantization is not straightforward because KV tensors require the input to be accessed. To address this, we employ a calibration dataset (a significantly smaller set of input samples compared to the training dataset) to collect the necessary tensors. Once the full-precision KV tensors at the $k$-th layer, $\mathbf{K}^{k}_{F}$ and $\mathbf{V}^{k}_{F}$, are obtained, they can be quantized using the same method to produce $\mathbf{K}^{k}_{Q}$ and $\mathbf{V}^{k}_{Q}$ based on Eqn. 3.7.
@@ 4 Experiments
We detail the experimental setup and model configurations in Sec. 4.1, followed by an analysis of the quantitative and qualitative performance of our approach in Sec. 4.2 and Sec. 4.4. Finally, we ablate the effectiveness of each component in our model in Sec. 4.3.
@@ 4.1 Experimental Setup
4.1.1 Model Settings
We use the Video-LLaVA [Lin et al., 2024] framework to explore Video-LLM interpolation. [^1] We develop our method based on the Video-LLaVA codebase. The evaluation of our approach is conducted using the IG-VLM [Kim et al., 2024] codebase. We employ Vicuna-7B v1.5 as the LLM backbone, mirroring the architecture used in Video-LLaVA. The visual encoders are adapted from LanguageBind, and the text tokenization is handled by the LLaMA tokenizer, which includes approximately 32,000 classes. The projection layers consist of 2 fully connected layers, facilitating the integration of visual and textual data.
4.1.2 Data and Training Details
Consistent with our focus on developing a training-free method, our approach requires no traditional data training processes. This aspect underscores the innovative nature of our interpolation method, which leverages existing model architectures and tools without the need for retraining or additional data, simplifying the implementation and reducing the computational overhead typically associated with training new models.
4.1.3 Datasets and Evaluation Metric
We assess INTP across a variety of zero-shot video question-answer benchmarks, which we classify as either open-ended or multiple-choice, depending on the type of question posed [Kim et al., 2024]. Our evaluation for open-ended VQA includes MSVD-QA [Xu et al., 2017], MSRVTT-QA [Xu et al., 2017], and ActivityNet-QA [Yu et al., 2019] benchmarks. We utilize GPT-assisted assessments as outlined in Video-ChatGPT [Maaz et al., 2023], which provide a thorough evaluation of the model’s response accuracy and correctness. For multiple-choice VQA tasks, we evaluate the model’s performance on NExT-QA [Xiao et al., 2021] and EgoSchema [Mangalam et al., 2024]. We determine accuracy by the model’s ability to correctly select the appropriate answer from the provided options.
Table 2: A comparison of different LVMs on video reasoning benchmarks. Like Video-LLaVA [Lin et al., 2024], ChatGPT-Assistant is employed to evaluate the following performance. The version of ChatGPT is “gpt-3.5-turbo”.
@@ 4.1.3 Datasets and Evaluation Metric
FrozenBiLM [Yang et al., 2022] 1B 32.2 - 16.8 - 24.7 -
VideoChat [Li et al., 2023a] 7B 56.3 2.8 45.0 2.5 - 2.2
LLaMA-Adapter [Zhang et al., 2023b] 7B 54.9 3.1 43.8 2.7 34.2 2.7
Video-LLaMA [Zhang et al., 2023a] 7B 51.6 2.5 29.6 1.8 12.4 1.1
Video-ChatGPT [Maaz et al., 2023] 7B 64.9 3.3 49.3 2.8 35.2 2.7
Video-LLaVA [Lin et al., 2024] 7B 70.7 3.9 59.2 3.5 45.3 3.3
@@
FrozenBiLM [Yang et al., 2022] 1B 26.9 - - - -
InternVideo [Wang et al., 2024] 1.3B 32.1 48.0 43.4 65.1 59.1
Sevilla [Yu et al., 2024] 7B - 61.3 61.5 75.6 63.6
Video-ChatGPT [Maaz et al., 2023] 7B - 61.9 57.4 69.9 61.7
Video-LLaVA [Lin et al., 2024] 7B 37.0 61.2 54.2 71.1 60.5
@@ 4.3 Ablation Studies
Figure 3: Video content and questions from ActivityNet [Yu et al., 2019]. The standard Video-LLaVA model exhibits limitations in accurately answering the questions, primarily due to its inability to process a sufficient number of frames (only the pink labeled frames are fed into Video-LLaVA). This constraint significantly hinders its effectiveness in managing complex video question-answering tasks. With our proposed INTP, the enhanced Video-LLM can process extended sequences of video frames. This not only addresses the frame limitations but also substantially enhances the model’s understanding in complex video question-answering scenarios.
@@ 4.4 Qualitative Results
Finally, we visualize the question-and-answer samples within the ActivityNet dataset to assess the qualitative performance of our approach in Fig. 3. In the first example, while Video-LLaVA incorrectly describes a person “riding a bike on a skateboard” and misidentifies the hair length, INTP-Video-LLaVA accurately describes the scene as a person “wearing orange clothes and jumping on stilts on a trampoline” and correctly states that the person does not have long hair. This shows INTP-Video-LLaVA’s superior ability to process and interpret a larger number of video frames, leading to more accurate scene description and detail recognition. In the second example, Video-LLaVA hallucinates the presence of a second person playing a guitar, which is not shown in the video frames. In contrast, INTP-Video-LLaVA correctly identifies that there are two people in the video, accurately describes the main activity (playing drums), and even provides specific details about the types of drums being played (congas and tabla). Additionally, INTP-Video-LLaVA correctly identifies the gender of the person playing the drums, demonstrating its improved ability to process and understand visual information across a longer sequence of frames. These examples highlight how INTP-Video-LLaVA’s capacity to handle more video frames leads to more accurate, detailed, and coherent descriptions of video content, significantly reducing errors and hallucinations compared to the standard Video-LLaVA model.
@@ 5 Conclusion
In conclusion, we presented INTP, an innovative interpolation method for Video-LLMs that effectively addresses the limitations of current models in processing extended video sequences. By developing an alternative video token rearrangement technique and a training-free LLM context window extension, INTP enables Video-LLMs to handle significantly more visual data (32 frames) without the need for additional training or substantial computational resources. Furthermore, the introduction of a KV-cache compression module optimizes memory usage during inference, enhancing deployment efficiency. These advancements not only extend the practical utility of Video-LLMs but also demonstrate the potential of training-free approaches in advancing Video-LLMs.
Societal Impacts. Our approach can democratize the access to advanced video processing technologies by reducing computational demands. Despite these advancements, it is important to note that INTP does not address potential security concerns related to the misuse of Video-LLMs. The ease of accessibility and enhanced capabilities could potentially be exploited by malicious actors to generate or manipulate video content in harmful ways. Thus, while INTP expands the practical applications of Video-LLMs and reduces barriers to their use, it also necessitates careful consideration of the ethical implications and security measures associated with their deployment.
Acknowledgement We thanks for the help from Yang Zhou, Haoxuan Wang, Zhihang Yuan for discussing the long-content LLM and KV-cache compression.